# Libo Qin  Qiguang Chen  Xiachong Feng  Yang Wu  Yongheng Zhang Yinghui Li  Min Li  Wanxiang Che  Philip S. Yu

### v1

@@ Abstract
While large language models (LLMs) like ChatGPT have shown impressive capabilities in Natural Language Processing (NLP) tasks, a systematic investigation of their potential in this field remains largely unexplored. This study aims to address this gap by exploring the following questions: (1) How are LLMs currently applied to NLP tasks in the literature? (2) Have traditional NLP tasks already been solved with LLMs? (3) What is the future of the LLMs for NLP? To answer these questions, we take the first step to provide a comprehensive overview of LLMs in NLP. Specifically, we first introduce a unified taxonomy including (1) parameter-frozen application and (2) parameter-tuning application to offer a unified perspective for understanding the current progress of LLMs in NLP. Furthermore, we summarize the new frontiers and the associated challenges, aiming to inspire further groundbreaking advancements. We hope this work offers valuable insights into the potential and limitations of LLMs in NLP, while also serving as a practical guide for building effective LLMs in NLP.
Large Language Models Meet NLP: A Survey
Libo Qin  Qiguang Chen  Xiachong Feng  Yang Wu  Yongheng Zhang Yinghui Li  Min Li  Wanxiang Che  Philip S. Yu Central South University   Harbin Institute of Technology   University of Hong Kong Tsinghua University   University of Illinons at Chicago lbqin@csu.edu.cn, {qgchen,car}@ir.hit.edu.cn
@@ 1 Introduction
Recently, large language models (LLMs) represent a significant breakthrough in AI through scaling up language models Zhao et al. (2023a); Kaddour et al. (2023); Yang et al. ; Hadi et al. (2023); Zhuang et al. (2023). Current studies on LLMs, such as GPT-series Brown et al. (2020); Ouyang et al. (2022), PaLM-series Chowdhery et al. (2022), OPT Zhang et al. (2022a), and LLaMA Touvron et al. (2023), have shown impressive zero-shot performance. In addition, LLMs also bring some emergent abilities including instruction following Wei et al. (2022a), chain-of-thought reasoning Wei et al. (2022c) and in-context learning Min et al. (2022), which attract increasing attention Wei et al. (2022b).
With the advancement of large language models, as shown in Figure 1, LLMs allow various natural language processing (NLP) tasks (e.g., zero-shot mathematical reasoning, text summarization, machine translation, information extraction and sentiment analysis) to be achieved through a unified generative paradigm, which has achieved remarkable success Wei et al. (2022c, 2023a); Qin et al. (2023a); Wang et al. (2023a, d, h, j); Wan et al. (2023b); Peng et al. (2023); Huang et al. (2023a). Additionally, some LLMs in NLP work without needing any additional training data and can even surpass traditional models fine-tuned with supervised learning. This advancement significantly contributes to the development of NLP literature. As a result, the community has witnessed an exponential growth of LLMs for NLP studies, which motivates us to investigate the following questions: (1) How are LLMs currently applied to NLP tasks in the literature? (2) Have traditional NLP tasks already been solved with LLMs? (3) What is the future of the LLMs for NLP?
@@
To answer the above questions, we make the first attempt to present a comprehensive and detailed analysis on LLMs for NLP. The overarching goal of this work is to explore the current developments in LLMs for NLP. To this end, in this paper, we first introduce the relevant background and preliminary. Furthermore, we introduce a unified paradigm on LLMs for NLP: (1) parameter-frozen application including (i) zero-shot learning and (ii) few-shot learning; (2) parameter-tuning application containing (i) full-parameter tuning and (ii) parameter-efficient tuning, aiming to provide a unified perspective to understand the current progress of LLMs for NLP:
  • Parameter-frozen application directly applies prompting approach on LLM for NLP tasks without the need for parameter tuning. This category includes zero-shot and few-shot learning, depending on whether the few-shot demonstrations is required.
  • Parameter-tuning application refers to the need for tuning parameters of LLMs for NLP tasks. This category includes both full-parameter and parameter-efficient tuning, depending on whether fine-tuning is required for all model parameters.
Finally, we conclude by identifying potential frontier areas for future research, along with the associated challenges to stimulate further exploration.
In summary, this work offers the following contributions:
@@
    New taxonomy: We introduce a new taxonomy including (1) parameter-frozen application and (2) parameter-tuning application, which provides a unified view to understand LLMs for NLP tasks.
@@
We expect this work will be a valuable resource for researchers and spur further advancements in the field of LLM-based NLP.
@@ 2.1 Parameter-Frozen Paradigm
Parameter-frozen paradigm can directly apply prompting for NLP tasks without any parameter tuning. As shown in Figure 2 (a), this category encompasses zero-shot learning and few-shot learning Brown et al. (2020); Kojima et al. (2022).
@@ 2.2 Parameter-Tuning Paradigm
^$\hat{\mathcal{M}}=\texttt{Fine-tune}(\mathcal{M}|\mathcal{D}),$ (3)
@@
Parameter-efficient tuning (PET) involves adjusting a set of existing parameters or incorporating additional tunable parameters (like Bottleneck Adapter Houlsby et al. (2019), Low-Rank Adaptation (LoRA) Hu et al. (2021), Prefix-tuning Li and Liang (2021a), and QLoRA Dettmers et al. (2023)) to efficiently adapt models for specific NLP tasks. Formally, parameter-efficient tuning first tunes a set of parameters $\mathcal{W}$, denoting as:
@@ 3 Natural Language Understanding
As shown in Figure 3, we first describe some typical NLP understanding tasks, which consists of Semantic Analysis (§$\S$3.1), Information Extraction (§$\S$3.2), Dialogue Understanding (§$\S$3.3), and Table Understanding (§$\S$3.4).
@@ 3.1 Sentiment Analysis
Sentiment analysis, a key function in natural language processing, identifies the emotional tone of a text, like positive opinions or criticisms Wankhade et al. (2022).
@@ 3.1.1 Parameter-Frozen Paradigm
With the help of instruction tuning, LLMs have been equipped with excellent zero-shot learning ability Belkhir and Sadat (2023). Recent studies Zhang et al. (2023g) find that using simple instructions can elicit ChatGPT’s strong capabilities on a series of sentiment analysis tasks such as sentiment classification and aspect-based sentiment analysis. Moreover, current mainstream LLMs Koto et al. (2024) possess the ability of multilingual understanding to analyze the sentiment conveyed by different languages based on sentiment lexicons Koto et al. (2024).
@@
Few-shot prompting not only elicits in-context learning in LLMs but also elaborates the intent of users more clearly. According to the findings presented by previous studies Zhang et al. (2023g); Zhao et al. (2023b); Xu et al. (2023c), incorporating exemplars to the prompts significantly boosts LLMs’ performance on aspect-based sentiment analysis and emotion recognition tasks. Furthermore, Sun et al. (2023b) introduce few-shot learning on more complex procedures, incorporating multi-LLM negotiation framework for sentiment analysis.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8.4em,font=,, where level=2text width=11.3em,font=,, where level=3text width=8.0em,font=,, where level=4text width=12em,font=,, [Parameter-Frozen Paradigm Taxonomy,ver [  Understanding (§3),ver [Sentiment Analysis (§3.1) [e.g., Zhang et al. (2023g), Koto et al. (2024), Zhao et al. (2023b), Xu et al. (2023c), Sun et al. (2023b) , leaf, text width=48.3em] ] [Information Extraction
 (§3.2) [e.g., Zhang et al. (2023c),Wei et al. (2023a),Xie et al. (2023), Li and Zhang (2023),Li et al. (2023e), Bi et al. (2023) , leaf, text width=48.3em] ] [Dialogue Understanding
 (§3.3) [e.g., Pan et al. (2023), He and Garner (2023), Hudeček and Dušek (2023), Heck et al. (2023), Gao et al. (2023a),
Li et al. (2022b), Zhang et al. (2023i, h), Wu et al. (2023c), Das et al. (2023a), Chi et al. (2023), Hu et al. (2022b),
King and Flanigan (2023), Addlesee et al. (2023),Chung et al. (2023),Lee et al. (2023),Lin et al. (2023), Cao (2023) , leaf, text width=48.3em] ] [Table Understanding (§3.4) [e.g., Singha et al. (2023), Patnaik et al. (2024), Ye et al. (2023, 2024), Sui et al. (2023a, b), Cheng et al. (2022),
Zhang et al. (2023f, k, j, 2024), Chen (2023), Luo et al. (2023b), Li et al. (2023b), Jiang et al. (2023) , leaf, text width=48.3em] ] ] [      Generation (§4),ver [Summarization (§4.1) [e.g., Goyal et al. (2022), Ravaut et al. (2023b), Bhaskar et al. (2022), Wang et al. (2023b), Zhang et al. (2023e, b),
Adams et al. (2023), Tang et al. (2023) , leaf, text width=48.3em] ] [Code Generation (§4.2) [e.g., Chen et al. (2021), Nijkamp et al. (2022), Christopoulou et al. (2022), Luo et al. (2023c), Allal et al. (2023),
Li et al. (2023f, g), Guo et al. (2024) Roziere et al. (2023), Zheng et al. (2023), , leaf, text width=48.3em] ] [Machine Translation (§4.3) [e.g., Wei et al. (2023b),Zhu et al. (2023a),Li et al. (2023a, c),Alves et al. (2023),Raunak et al. (2023),Lu et al. (2023b) , leaf, text width=48.3em] ] [Mathematical Reasoning
 (§4.4) [e.g., Wei et al. (2022c), Zhang et al. (2022b), Kojima et al. (2022), Wang et al. (2023g), Touvron et al. (2023),
Lu et al. (2023d), Gao et al. (2023b) , leaf, text width=48.3em] ]] ]
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8.4em,font=,, where level=2text width=11.3em,font=,, where level=3text width=8.0em,font=,, where level=4text width=12em,font=,, [Parameter-Tuning Paradigm Taxonomy,ver [  Understanding (§3),ver [Sentiment Analysis (§3.1) [e.g., Wang et al. (2022), Varia et al. (2022), Yang and Li (2023), Zhao et al. (2016), Qiu et al. (2023) , leaf2, text width=48.3em] ] [Information Extraction
 (§3.2) [e.g., Lu et al. (2023a),Gan et al. (2023),Sainz et al. (2023),Wang et al. (2023f), Das et al. (2023b),Liang et al. (2023) , leaf2, text width=48.3em] ] [Dialogue Understanding
 (§3.3) [e.g., Xie et al. (2022),Zhao et al. (2022a), Gupta et al. (2022),Yu et al. (2022),Feng et al. (2023b), Liu et al. (2023b) , leaf2, text width=48.3em] ] [Table Understanding (§3.4) [e.g., Li et al. (2023d), Xie et al. (2022), Xue et al. (2023), Zhang et al. (2023a), Zhu et al. (2024), Bai et al. (2023),
Zhang et al. (2023d) , leaf2, text width=48.3em] ]] [       Generation (§4),ver [Summarization (§4.1) [e.g., Pagnoni et al. (2022), Zhao et al. (2022b), Yuan et al. (2022), Feng et al. (2023a), Li and Liang (2021b),
Ravaut et al. (2023a) , leaf2, text width=48.3em] ] [Code Generation (§4.2) [e.g., Wang et al. (2021, 2023i),Le et al. (2022),Shojaee et al. (2023),Ayupov and Chirkova (2022),Zhuo et al. (2024),
Weyssow et al. (2023) , leaf2, text width=48.3em] ] [Machine Translation (§4.3) [e.g., Xu et al. (2023b, 2024), Iyer et al. (2023),Moslem et al. (2023),Ustun and Stickland (2022),Alves et al. (2023)
Wu et al. (2023a, 2024), , leaf2, text width=48.3em] ] [Mathematical Reasoning
 (§4.4) [e.g.,  Luo et al. (2023a), Yue et al. (2023), Ho et al. (2023), Schick et al. (2023), Hu et al. (2022a, 2023b) , leaf2, text width=48.3em] ] ] ]

### v2

# Large Language Models Meet NLP: A Survey

@@ Abstract
While large language models (LLMs) like ChatGPT have shown impressive capabilities in Natural Language Processing (NLP) tasks, a systematic investigation of their potential in this field remains largely unexplored. This study aims to address this gap by exploring the following questions: (1) How are LLMs currently applied to NLP tasks in the literature? (2) Have traditional NLP tasks already been solved with LLMs? (3) What is the future of the LLMs for NLP? To answer these questions, we take the first step to provide a comprehensive overview of LLMs in NLP. Specifically, we first introduce a unified taxonomy including (1) parameter-frozen paradigm and (2) parameter-tuning paradigm to offer a unified perspective for understanding the current progress of LLMs in NLP. Furthermore, we summarize the new frontiers and the corresponding challenges, aiming to inspire further groundbreaking advancements. We hope this work offers valuable insights into the potential and limitations of LLMs, while also serving as a practical guide for building effective LLMs in NLP.
keywords:
Natural Language Processing, Large Language Models, Parameter-frozen Paradigm, Parameter-tuning Paradigm, ChatGPT
[…]
@@ 1 Introduction
Recently, large language models (LLMs) represent a significant breakthrough in AI through scaling up language models [1, 2, 3, 4, 5, 6, 7, 8, 9]. Current studies on LLMs, such as GPT-series [10, 11], PaLM-series [12], OPT [13], and LLaMA [14], have shown impressive zero-shot performance. In addition, LLMs also bring some emergent abilities including instruction following [15], chain-of-thought reasoning [16] and in-context learning [17], which attract increasing attention [18].
With the advancement of large language models, as shown in Figure 1, LLMs allow various natural language processing (NLP) tasks (e.g., zero-shot mathematical reasoning [16, 19], text summarization [20, 21], machine translation [22, 23], information extraction [24, 25] and sentiment analysis [26, 27]) to be achieved through a unified generative paradigm, which has achieved remarkable success [1, 28, 29, 30]. Additionally, some LLMs in NLP work without needing any additional training data and can even surpass traditional models fine-tuned with supervised learning. This advancement significantly contributes to the development of NLP. As a result, the community has witnessed an exponential growth of LLMs for NLP studies, which motivates us to investigate the following questions: (1) How are LLMs currently applied to NLP tasks in the literature? (2) Have traditional NLP tasks already been solved with LLMs? (3) What is the future of the LLMs for NLP?
@@
To answer the above questions, we present a comprehensive and detailed analysis on LLMs from the perspective of independent NLP tasks. The overarching goal of this work is to explore current developments in LLMs for NLP. To this end, in this paper, we first introduce the relevant background and preliminary. Furthermore, we introduce a unified paradigm on LLMs for NLP: (1) parameter-frozen paradigm including (i) zero-shot learning and (ii) few-shot learning; (2) parameter-tuning paradigm containing (i) full-parameter tuning and (ii) parameter-efficient tuning, aiming to provide a unified perspective to understand the current progress of LLMs for NLP tasks:
  • Parameter-frozen paradigm directly applies prompting approach on LLM for NLP tasks without the need for parameter tuning. This category includes zero-shot and few-shot learning, depending on whether the few-shot demonstrations is required.
  • Parameter-tuning paradigm refers to the need for tuning parameters of LLMs for NLP tasks. This category includes both full-parameter and parameter-efficient tuning, depending on whether fine-tuning is required for all model parameters.
Finally, we conclude by identifying potential frontier areas for future research, along with the associated challenges to stimulate further exploration. In summary, this work offers the following contributions:
@@
    New taxonomy: We introduce a new taxonomy including (1) parameter-frozen paradigm and (2) parameter-tuning paradigm, which provides a unified view to understand LLMs for NLP tasks.
@@
We hope this work will be a valuable resource for researchers and spur further advancements in the field of LLM-based NLP.
@@ 2.1 Parameter-Frozen Paradigm
Parameter-frozen paradigm can directly apply prompting for NLP tasks without any parameter tuning. As shown in Figure 2 (a), this category encompasses zero-shot learning and few-shot learning [10, 31].
@@ 2.2 Parameter-Tuning Paradigm
$\hat{\mathcal{M}}=\texttt{Fine-tune}(\mathcal{M}|\mathcal{D}),$ (3)
@@
Parameter-efficient tuning (PET) involves adjusting a set of existing parameters or incorporating additional tunable parameters (like Bottleneck Adapter [32], Low-Rank Adaptation (LoRA) [33], Prefix-tuning [34], and QLoRA [35]) to efficiently adapt models for specific NLP tasks. Formally, parameter-efficient tuning first tunes a set of parameters $\mathcal{W}$, denoting as:
@@
Table 1: Comparison of resource consumption and performance across NLP adaptation paradigms. Data is compiled from Mundra et al. [36], Hu et al. [37], Dettmers et al. [35].[+]: better performance / low resource consumption, [++]: much better performance / moderate resource consumption, [+++]: best performance / high resource consumption, [-]: no consumption. Zero-shot Learning has lowest resource consumption and best out-of-domain task generalization, while Full-Parameter Tuning has highest cost and best in-domain performance.
Strategy Training Cost Memory (Train) Memory (Infer) Latency Accuracy Generalization
Zero-Shot Learning \$0 - + + + +++
Few-Shot Learning \$0 - ++ ++ ++ ++
Full Parameter-Tuning ¿\$1K 2 × model size + + +++ +
[…]
2.3 Comparison of Paradigms
To further understand the advantages on different paradigms, we summarize the resource consumption and performance of each paradigm in Table 1. Generally speaking, zero-shot learning offers the highest application efficiency, moderate improvements on in-domain tasks, and robust out-of-domain generalization. In contrast, few-shot learning typically yields superior in-domain performance relative to zero-shot learning; however, it demands greater computational resources, achieves lower overall efficiency, and exhibits reduced generalization to novel domains. Full-parameter tuning, when ample training data and resources are available, attains the best in-domain performance but at the expense of the least efficient deployment and the weakest transfer to out-of-domain settings. Finally, parameter-efficient tuning strikes a balance: with limited resources, it can match or exceed the performance of full-parameter tuning in certain cases, while offering higher efficiency and often improved generalization beyond the training domain.
@@ 3 Natural Language Understanding
As shown in Figure 3, we first describe some typical NLP understanding tasks, which consists of Semantic Analysis ($\S$3.1), Information Extraction ($\S$3.2), Dialogue Understanding ($\S$3.3), and Table Understanding ($\S$3.4).
@@ 3.1 Sentiment Analysis
Sentiment analysis, a key function in natural language processing, identifies the emotional tone of a text, like positive opinions or criticisms [38].
@@ 3.1.1 Parameter-Frozen Paradigm
With the help of instruction tuning, LLMs have been equipped with excellent zero-shot learning ability [39]. Recent studies [40] find that using simple instructions can elicit ChatGPT’s strong capabilities on a series of sentiment analysis tasks such as sentiment classification and aspect-based sentiment analysis. Current mainstream LLMs possess the ability of multilingual understanding to analyze the sentiment conveyed by different languages based on sentiment lexicons [41]. Moreover, Du et al. [42] propose a prompting framework to evaluate and reveal LLMs’ limitations in financial attribute reasoning for sentiment analysis, highlighting weaknesses in numerical and understanding.
@@
Few-shot prompting not only elicits in-context learning in LLMs but also elaborates the intent of users more clearly. According to the findings presented by previous studies [40, 43, 44, 45], incorporating exemplars to the prompts significantly boosts LLMs’ performance on aspect-based sentiment analysis and emotion recognition tasks. Furthermore, Sun et al. [46] introduce few-shot learning on more complex procedures, incorporating multi-LLM negotiation framework for deeper sentiment analysis.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8.4em,font=,, where level=2text width=11.3em,font=,, where level=3text width=8.0em,font=,, where level=4text width=12em,font=,, [Parameter-Frozen Paradigm Taxonomy,ver [ Understanding (§3),ver [Sentiment Analysis (§3.1) [e.g., Zhang et al. [40], Koto et al. [41], Zhao et al. [43], Xu et al. [44], Sun et al. [46], Du et al. [42], Zhang et al. [47] , leaf, text width=48.3em] ] [Information Extraction
 (§3.2) [e.g., Zhang et al. [48],Wei et al. [24],Xie et al. [49], Li and Zhang [50],Li et al. [51], Bi et al. [52], Fornasiere et al. [53],
Tang et al. [54] , leaf, text width=48.3em] ] [Dialogue Understanding
 (§3.3) [e.g., Pan et al. [55], He and Garner [56], Hudeček and Dušek [57], Heck et al. [58], Gao et al. [59], Li et al. [60],
Zhang et al. [61, 62], Wu et al. [63], Das et al. [64], Chi et al. [65], Hu et al. [66], King and Flanigan [67], Addlesee et al. [68],
Chung et al. [69], Lee et al. [70], Lin et al. [71], Cao [72] , leaf, text width=48.3em] ] [Table Understanding (§3.4) [e.g., Singha et al. [73], Patnaik et al. [74], Ye et al. [75, 76], Sui et al. [77, 78], Cheng et al. [79], Zhang et al. [80, 81, 82, 83],
Chen [84], Luo et al. [85], Li et al. [86], Jiang et al. [87], Wang et al. [88], Kong et al. [89] , leaf, text width=48.3em] ] ] [   Generation (§4),ver [Summarization (§4.1) [e.g., Goyal et al. [90], Ravaut et al. [91], Bhaskar et al. [92], Wang et al. [93], Zhang et al. [94, 95], Adams et al. [96], Tang et al. [97] , leaf, text width=48.3em] ] [Code Generation (§4.2) [e.g., Chen et al. [98], Nijkamp et al. [99], Christopoulou et al. [100], Luo et al. [101], Allal et al. [102], Li et al. [103, 104],
Guo et al. [105], Roziere et al. [106], Zheng et al. [107], , leaf, text width=48.3em] ] [Machine Translation (§4.3) [e.g., Wei et al. [108],Zhu et al. [109],Li et al. [110, 111],Alves et al. [112],Raunak et al. [113],Lu et al. [114] , leaf, text width=48.3em] ] [Mathematical Reasoning
 (§4.4) [e.g., Wei et al. [16], Zhang et al. [115], Kojima et al. [31], Wang et al. [116], Touvron et al. [14], Lu et al. [117], Gao et al. [118],
Das et al. [119] , leaf, text width=48.3em] ]] ]
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8.4em,font=,, where level=2text width=11.3em,font=,, where level=3text width=8.0em,font=,, where level=4text width=12em,font=,, [Parameter-Tuning Paradigm Taxonomy,ver [ Understanding (§3),ver [Sentiment Analysis (§3.1) [e.g., Wang et al. [120], Varia et al. [121], Yang and Li [122], Zhao et al. [123], Qiu et al. [124] , leaf2, text width=48.3em] ] [Information Extraction
 (§3.2) [e.g., Lu et al. [125],Gan et al. [126],Sainz et al. [127],Wang et al. [128], Das et al. [129], Liang et al. [130], Dagdelen et al. [131],
Xue et al. [132], Rixewa et al. [133] , leaf2, text width=48.3em] ] [Dialogue Understanding
 (§3.3) [e.g., Xie et al. [134],Zhao et al. [135], Gupta et al. [136],Yu et al. [137],Feng et al. [138], Liu et al. [139] , leaf2, text width=48.3em] ] [Table Understanding (§3.4) [e.g., Li et al. [140], Xie et al. [134], Xue et al. [141], Zhang et al. [142], Zhu et al. [143], Bai et al. [144], Zhang et al. [145],
He et al. [146], Li et al. [147] , leaf2, text width=48.3em] ]] [    Generation (§4),ver [Summarization (§4.1) [e.g., Pagnoni et al. [148], Zhao et al. [149], Yuan et al. [150], Feng et al. [151], Li and Liang [152], Ravaut et al. [153] , leaf2, text width=48.3em] ] [Code Generation (§4.2) [e.g., Wang et al. [154, 155],Le et al. [156],Shojaee et al. [157],Ayupov and Chirkova [158],Zhuo et al. [159], Weyssow et al. [160] , leaf2, text width=48.3em] ] [Machine Translation (§4.3) [e.g., Xu et al. [161, 162], Iyer et al. [163],Moslem et al. [164],Ustun and Stickland [165],Alves et al. [112] Wu et al. [166, 167], , leaf2, text width=48.3em] ] [Mathematical Reasoning
 (§4.4) [e.g.,  Luo et al. [168], Yue et al. [169], Ho et al. [170], Schick et al. [171], Hu et al. [37, 172], Shi et al. [173], Shao et al. [174],
Luo et al. [175], Chen et al. [176], Liu et al. [177], Yu et al. [178], Ranaldi et al. [179], Srivastava et al. [180], Cai et al. [181] , leaf2, text width=48.3em] ] ] ]